OpenStudy (anonymous):

7x+4y=-4 5x+8y=28 elimination/solve. multiplacation or addition subtraction.

OpenStudy (mertsj):

Multiply the top equation by -2. Add the result to the bottom equation.

OpenStudy (anonymous):

14x+8y=-8?

OpenStudy (anonymous):

..............

OpenStudy (phoenixfire):

-2. not 2. You want to eliminate the 'y' term. then you solve the resulting equation for x. and substitute that value back into either of the first equations to get a value for y.

OpenStudy (anonymous):

9x=20

OpenStudy (anonymous):

!

OpenStudy (anonymous):

mersj...

OpenStudy (phoenixfire):

7x+4y=-4 (1) 5x+8y=28 (2) -14x-8y=8 <= (1) * -2 = (3) -9x=36 <= (3) + (2) Solve for x: x=-4 Now plug -4 into either (1) or (2) and solve for y.

OpenStudy (anonymous):

.....i have no words to express my gratitude.